OpenStudy (anonymous):

Solve: E = 6.625 X 10- X 7.309 X 10 1sec

OpenStudy (anonymous):

energy = E = constant, h Frequency, f

OpenStudy (anonymous):

could you write in other form your task form Solve: to the end cause this is confusing what you wrote

OpenStudy (anonymous):

thts all it gave me in the book

OpenStudy (anonymous):

but write it more properly when its written like this i cant see what is what

OpenStudy (anonymous):

il try

OpenStudy (anonymous):

could you write something more like this: E= t= and so on

OpenStudy (anonymous):

it's actually easier if you try to draw it. are you trying to find frequency, energy, or wavelength. those are the only possibilities with this info

OpenStudy (anonymous):

isnt h = 6.625*10^-34 and does ur question say the frequency is 7.309 inverse seconds?

OpenStudy (anonymous):

thats the problem i cant read what is what

OpenStudy (anonymous):

me nether

OpenStudy (anonymous):

what are you solving for. start with that. and then draw out exactly what it says in ur book

OpenStudy (anonymous):

wavelength

OpenStudy (anonymous):

Energy = (h)(speed of light)/ wavelength

OpenStudy (anonymous):

speed of light = (frequency)(wavelength) those are the two formulas that i use for the questions

OpenStudy (anonymous):

E = h * (ni - frequency) and c=(lambda - wavelenght) * (ni)

OpenStudy (anonymous):

oh i got it thanks for the help tho